OpenStudy (anonymous):

Does y =9? 4x + 3y = -1 3x + y = 3

OpenStudy (uri):

How did you do it? Show workl @aleximhappy08

OpenStudy (anonymous):

-8+3y=-1 --> 3y=8+1 --> 3y=9 -->y=3 -6+y=3 -->y=3+6 -->y=9 I'm not sure which is right I solved for x and got -2

OpenStudy (anonymous):

if you are using 9 for y then you plug the 9 in for y so it would be 4x+3(9)=-1

OpenStudy (anonymous):

then once you solve that you will get \[4x+3(9)=-1\] \[4x+27=-1\]

OpenStudy (anonymous):

x=6.5

OpenStudy (anonymous):

and then if you plug in 9 for the second equation x=-2

OpenStudy (anonymous):

I'm confuseed. If you use the elimination property to solve for x then x should equal -2

OpenStudy (texaschic101):

4x + 3y = -1 -->(3) 3x + y = 3-->(-4) --------------------- 12x + 9y = -3 -12x -4y = -12 ------------ 5y = -15 y = -3

OpenStudy (anonymous):

Ohhhhh

OpenStudy (anonymous):

Then x would equal? -2?

OpenStudy (texaschic101):

4x + 3y = -1 4x + 3(-3) = -1 4x -9 = -1 4x = -1 + 9 4x = 8 x = 2

OpenStudy (uri):

We put the value of x in the original equation (1) and solve for y.

OpenStudy (anonymous):

So its positive not negitive. Okay

OpenStudy (texaschic101):

check it.... 4x + 3y = -1 4(2) + 3(-3) = -1 8 - 9 = -1 -1 = -1 (correct) x = 2, y = -3

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

No actually thanks!

OpenStudy (texaschic101):

your welcome....you could have also solved by substitution

OpenStudy (anonymous):

true....